# Going beyond the creative limits of next-token prediction

### v3

@@ 3 Training and Inference
In our experiments, we use the score entropy discrete diffusion model (SEDD, Lou et al., 2023). Loosely, we can think of the objective here as a hybrid of the teacherless objective, where the input is fully masked or corrupted (and the model must unmask or uncorrupt all of it), and various easier objectives with partial maskings or corruptions of the input. During inference, the model starts with a fully masked or corrupted sequence of tokens, and iteratively corrects subsets of them.
Inference. In all the above techniques, we extract each sample independently from the model (as against say, extracting them in continuous succession in the same context). For Transformers, during inference, we perform standard autoregression in both the next- and multi-token trained settings. We do this either with greedy decoding or with nucleus sampling (Finlayson et al., 2024).
@@ 3.1 Seed-conditioning
Unlike closed-ended tasks where a prompt (i.e., a prefix) maps to a unique correct answer, our open-ended tasks are prompt-free. As an alternative to this, we also tried prepending a prompt of pause tokens (Goyal et al., 2024) to all datapoints — both during training and during inference — to enable extra computation before the model emits its outputs. Next, we tried a more sophisticated alternative we term seed-conditioning. Here, we associate each training point with an arbitrary (meaningless) prefix to the model (e.g., of random characters) called a seed; during test-time, we prompt with novel seeds hoping to extract novel test data. We provide hypotheses motivating this procedure in §5.1.
@@ 4 Experimental results
Part of our experiments are performed for a Gemma v1 (2B) pre-trained model (Gemma Team et al., 2024), averaged over 4 runs. For diffusion, we use a 90M (non-embedding) parameters Score Entropy Discrete Diffusion model (SEDD; Lou et al., 2023). For a fair comparison against NTP, we use a 86M (non-embedding) parameters GPT-2 model (Radford et al., 2019). In all our experiments, we finetune the models until it is clear that algorithmic creativity (Eq. 1) has saturated. All values are reported from this checkpoint. Finally, since our best Transformer results were under noise-conditioning (for both next- and mult-token training), our main results are reported under that training setting; we provide various ablations without that as well. Please see §D for more experimental details, and §C for precise dataset details (e.g., how the graph is constructed, how sequences are formatted etc.,).
@@ 4.1 Observations
Multi-token prediction improves algorithmic creativity significantly. In all our datasets, the algorithmic creativity of the Gemma v1 (2B) model increases significantly under multi-token prediction ( Fig 3), with nearly a  5x factor for the discovery datasets. Note that for this, we have selected the learning rate favorable towards next-token prediction; tuning for multi-token yields further gains (Fig 15).
@@
For the smaller models, in Fig 4, we find that the diffusion model improves over next-token training of similar-sized Transformers on all tasks, except on Sibling Discovery where it is mildly worse. With teacherless training, the gains are absent; however we did find that the creativity of the top-$K$ samples is generally improved (Fig 24), indicating a less peaky distribution. Regardless, this echoes prior discussions suggesting that for smaller Transformers, teacherless training is hard to optimize (B&N’24) and other multi-token objectives even hurt (Gloeckle et al., 2024) smaller models of the order of 300M.
Multi-token prediction reduces memorization for the larger model. Algorithmic creativity may suffer either because the model outputs incoherent garbage, or because it repeats the same original output, or because it simply parrots out the training data. For our larger model, it is the last reason that dominates: across the board (in Fig 3), next-token prediction is significantly prone to memorizing the data, while the multi-token method is highly resistant. As foreshadowed in §2.6, we hypothesize that this is because NTP memorizes the earlier training tokens without a global plan, having fit the later tokens via local coherence rules (because of Clever Hans cheats à la B&N’24). Note that the smaller models have much lower absolute memorization values; the multi-token objective preserves or mildly worsens this, in a way that does not always hurt creativity proportionally. We did find that amongst the top-$K$ samples, teacherless training has reduced memorization for some tasks (see Fig 25).
@@
Figure 6: Seed-conditioning achieves high algorithmic creativity in the GPT-2 (86M) model: All models are trained with NTP. The color denotes temperature, while the X-axis, the prefix. Remarkably, seed-conditioning achieves high creativity even with greedy decoding, and improves creativity on all except Sibling Discovery.
Seed-conditioning improves algorithmic creativity for Transformers. Orthogonal to the effect of multi-token vs. next-token objectives, we point out the effect of seed-conditioning a Transformer. First, and most remarkably, with seed-conditioning, there is no need for temperature: even deterministic generation via greedy decoding generates diverse outputs that have highly non-trivial creativity values (Fig 5, Fig 6). This means that, even though the seeds we provide are arbitrarily paired with the training datapoint, the models have learned to transform the arbitrary noise into meaningful randomness. We also find that increasing the seed lengths consistently boosts algorithmic creativity for both next-token and multi-token approaches (Fig 523). Thus, for Transformers, we propose viewing seed-conditioning as a knob for diversity distinct from temperature sampling.
Next, in comparison to temperature sampling, in all our settings, seed-conditioning results in the algorithmic creativity comparable to temperature sampling. In fact, for the larger models (Fig 5), seed-conditioning generally outperforms the baseline “null” conditioning (which appears mode-collapsed, see Fig 181920). For the smaller model as well (Fig 6), we find gains in all tasks except for Sibling Discovery. Besides, for any fixed temperature, prefixing a seed only improves performance over a null prefix (Fig 618).
@@
4.2 An initial exploration of real-world summarization
For a more realistic examination of our findings, we conduct preliminary investigation of GPT models finetuned with NTP and the multi-token teacherless objectives on summarization tasks (XSUM, CNN/DailyMail). We measure the diversity of a model for any given prompt by generating $5$ different completions and computing a Self-Bleu metric (Zhu et al., 2018). Admittedly though, a summarization task is not as open-ended as we would like: a higher quality model (i.e., higher Rouge; Lin, 2004) necessarily means lower diversity. To account for this, we plot how diversity evolves over time as a function of the quality of the model; we then find in Fig 7 that for a given model quality, the larger multi-token models achieve higher diversity (albeit only by a slight amount). This increase does not hold for smaller models and is not always noticeable for CNN/DailyMail (see §I). Interestingly, teacherless training consistently shows an increase in summarization quality, measured by Rouge.
@@ 4.2 An initial exploration of real-world summarization
Figure 7: Multi-token training improves diversity scores for XSUM summarization for large GPT-2 models: Here, we plot diversity and quality as measured over multiple checkpoints during finetuning, and observe differences in diversity for a fixed quality.
@@ 5.1 Intuition for seed-conditioning
Seed-conditioning produces non-trivial algorithmic creativity even with greedy decoding. Perhaps, one could view the seed prefixes as a simpler alternative to varying the wordings of a prompt (Li et al., 2023; Lau et al., 2024; Naik et al., 2024) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. However, it is unexpected that the seeds provide any meaningful randomness at all, given they are arbitrarily picked with no semantic relationship to the training point. We do not understand this presently.
@@ 7 Related Work
Most related to seed-conditioning is DeSalvo et al. (2024) who induce diversity by varying a soft-prompt learned using a reconstruction loss. Our approach requires no modification to the architecture or the loss; however, we train the whole model, which is more expensive. A concurrent position paper (Jahrens & Martinetz, 2025) conceptually suggests injecting noise with the same motivation as us. The benefits of seed-conditioning may also be related to the fact that varying the wording in a prompt is known to induce diverse outputs (Li et al., 2023; Lau et al., 2024; Naik et al., 2024). Various works also inject noise into a Transformer, in a different form (e.g., inducing Gaussian noise), and for a different function such as quality, robustness (Hua et al., 2022; Jain et al., 2024) or efficiency (Wang et al., 2024c).
@@ 8 Conclusions
This work provides a minimal test-bed of tasks abstracting distinct modes of creativity. While these tasks are admittedly an extreme caricaturization of real-world tasks, they enable us to quantify otherwise elusive metrics like originality and diversity. They also enable us to control and investigate distinct parts of the current apparatus for language modeling (next-token learning and softmax-based temperature sampling) and advocate for alternatives (multi-token learning and seed-conditioning). The surprising effectiveness of seed-conditioning raises various open questions (§B.1). There are also other profound questions as to whether reasoning-enhancing methods like RL and CoT are optimal for enhancing open-ended diversity and originality (§5.2). Overall, we hope our work inspires discussion in the various directions of multi-token prediction, creativity and planning.
@@ Appendix A Transformer Training Objectives
In teacherless (multi-token) training (Monea et al., 2023; Bachmann & Nagarajan, 2024; Tschannen et al., 2023), we make use of an uninformative input string $\bm{\$}$ that simply corresponds to a series of dummy tokens \$currency-dollar\\$\$.
@@ B.3 Examples of Triangle Discovery
    Antanaclasis: An antanaclasis involves using a word in two different senses in a sentence, while still ensuring that each sense has a coherent relationship with the rest of the sentence. Consider Benjamin Franklin’s quote, Your argument is sound, nothing but sound. Here, the two senses are $\texttt{sound}_{1}$ as in “logically correct”, $\texttt{sound}_{2}$ as in “noise”. This sentence encodes an pairwise relationship between three entities $\{\texttt{argument},\texttt{sound}_{1}$,$\texttt{sound}_{2}\}$ individually. While the last two entities (the two senses) themselves must be related to each other (through the common word, sound), for a coherent sentence, both senses must also be appropriate descriptors for the first entity, argument. Thus, constructing this sentence requires searching through one’s vocabulary to discover three words that satisfy these three relationships simultaneously.
@@
[…]
@@ C.1 Datasets inspired by combinational creativity
Next, we want a training dataset such that (a) the model can infer all the edges from the graph and yet (b) not all triangles appear in the dataset. This necessitates training on a dataset that consists not only of a subset of the triangles, but also of edges from the graph. Our training data consists of two parts: (1) $1/3$ are random triangles from the graph, (2) $2/3$ are random edges from the graph. In the training set, the triangle and edge samples are distinguished by a prefix “triangle:” or “edge:”. During test-time, we ensure that the model is prompted with “triangle:”. A triangle $(u,v,w)$ is tokenized as “tri: $(u,v),(v,w),(w,u)$” and an edge $(u,v)$ as “edge: $(u,v),(v,u)$”. We provide both the directions of edge to potentially avoid any issues with the reversal curse (Berglund et al., 2024; Allen-Zhu & Li, 2023a).
For the default setting of the dataset, we set $|{V}|=999$, $\texttt{deg}=3$, $\texttt{tri}=6$.
@@ Appendix D Further experimental details
Details for Gemma v1 (2B) model.  In Table 1, we provide the hyperparameter details for each of our datasets. We note some common details here. First, the batch size is $4$, but each sequence is packed with multiple examples; thus the model sequence length (divided by the input length) can be treated as a multiplicative factor that determines the effective batch size. The learning rates are chosen favorable to next-token prediction (not multi-token prediction). The training steps were chosen roughly based on a point after which the model had saturated in algorithmic creativity (and exhibited decreasing creativity). We use a learning rate with linear warm up for $100$ steps, followed by cosine annealing upto a factor 0.01×0.01$\times$0.01 × of the maximum learning rate. To measure creativity, we sample a test dataset $T$ of $1024$ datapoints.
We represent the main tokens in our tasks with integers (ranging upwards of 00 to as many distinct integers as required). In the seed-conditioning setting, we use seeds of default length $10$, using randomly sampled uppercase characters from the English alphabet. In all datasets, we space-separate the vertices in a string, and comma-separate the edges.
@@
Hyperparameter Sibling Discovery Triangle Discovery Circle Construction Line Construction
@@
Weight given to multi-token obj. $0.5$ $0.5$ $0.75$ $0.75$
@@
Hyperparameter Sibling Discovery Triangle Discovery Circle Construction Line Construction
Max. Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$ $1\times 10^{-4}$
@@
Weight given to multi-token obj. $0.5$ $0.33$ $0.75$ $0.75$
@@ E.1 Varying sampling methods
Fig 18, Fig 19, and Fig 20 report creativity, memorization and coherence (i.e., fraction of generated strings that are coherent) for various sampling methods (greedy decoding and nucleus sampling) with various prefix conditionings (namely, null, pause and seed).

### v4

@@ 3 Training and Inference
We use the score entropy discrete diffusion model (SEDD (90M), Lou et al., 2023). Loosely, we can think of the objective here as a hybrid of the teacherless objective, where the input is fully masked or corrupted (and the model must unmask or uncorrupt all of it), and various easier objectives with partial maskings or corruptions of the input. During inference, the model starts with a fully masked or corrupted sequence of tokens, and iteratively corrects subsets of them.
Inference. We extract independent samples from the model. For Transformers, during inference, we perform standard autoregression in both the next- and multi-token trained settings. We do this either with greedy decoding or with nucleus sampling (Finlayson et al., 2024).
@@ 3.1 Seed-conditioning
While temperature sampling is the standard way to elicit diversity from a Transformer, we speculate that this can lead to cognitive overload: the model must process multiple thoughts to compute a diverse softmax distribution. Alternatively, we propose conditioning on a random seed as input to the model, hoping that a thought can be fleshed out with singular focus. Thus, during training, each point is associated with an arbitrary (meaningless) prefix (e.g., of random characters) called a seed (note that there is no prefix in our tasks besides the seed). During test-time, novel seeds are used to extract test data. We can also view seed-conditioning more neutrally as a controlled way to simulate variations in prompt-wordings for a fixed task (e.g., design a word problem vs construct a problem). We elaborate on these intuitions in §5.1.
@@ 4 Experimental results
Part of our experiments are performed for a Gemma v1 (2B) pre-trained model (Gemma Team et al., 2024), averaged over 4 runs. For diffusion, we use a 90M (non-embedding) parameters Score Entropy Discrete Diffusion model (SEDD (90M); Lou et al., 2023). For a fair comparison against NTP, we use a 86M (non-embedding) parameters GPT-2 (86M) model (Radford et al., 2019). In all our experiments, we finetune the models until it is clear that algorithmic creativity (Eq. 1) has saturated. All values are reported from this checkpoint. Finally, since our best Transformer results were under seed-conditioning (for both next- and mult-token training), our main results are reported under that training setting; we provide various ablations without that as well. Please see §D for more experimental details, and §C for precise dataset details (e.g., how the graph is constructed, how sequences are formatted etc.,).
@@ 4.1 Observations
Multi-token prediction improves algorithmic creativity significantly. In all our datasets, the algorithmic creativity of the Gemma v1 (2B) model increases significantly under multi-token prediction ( Fig 3), with nearly a  5x factor for the discovery datasets. Note that for this, we have selected the learning rate favorable towards next-token prediction; tuning for multi-token yields further gains (Fig 15). For the smaller models, in Fig 4, we find that the diffusion model improves over next-token training of similar-sized Transformers on all tasks, except on Sibling Discovery where it is mildly worse. With teacherless training, the gains are absent; however we find that the creativity of the top-$K$ samples is generally improved (Fig 24), indicating a less peaky distribution. Regardless, the overall trend echoes prior findings that for smaller Transformers, teacherless training is hard to optimize and other multi-token objectives even hurt (Gloeckle et al., 2024) smaller models.
@@
Multi-token prediction reduces memorization for the larger model. Algorithmic creativity may suffer either due to either incoherence or mode collapse or memorization. For our larger model, it is the last reason that dominates: across the board (in Fig 3), next-token prediction is significantly prone to memorizing the data, while the multi-token method is highly resistant. As foreshadowed in §2.6, we hypothesize that this is because NTP memorizes the earlier training tokens without a global plan, having fit the later tokens via local coherence rules (because of Clever Hans cheats à la B&N’24). In contrast, the smaller models have much lower absolute memorization values; the multi-token objective preserves or mildly worsens this, in a way that does not always hurt creativity proportionally. here, we find that amongst the top-$K$ samples, teacherless training has reduced memorization for some tasks (see Fig 25).
@@
Figure 6: Seed-conditioning achieves comparable algorithmic creativity in the GPT-2 (86M) model: All models are trained with NTP. The color denotes temperature, while the X-axis, the prefix. Remarkably, seed-conditioning achieves comparable creativity even with greedy decoding, and improves creativity particularly in our exploratory creativity tasks (bottom).
Seed-conditioning improves algorithmic creativity for Transformers. Orthogonal to the effect of multi-token vs. next-token objectives, we point out the effect of seed-conditioning a Transformer. First, and most remarkably, with seed-conditioning, there is no need for temperature: even deterministic greedy decoding generates diverse outputs and highly non-trivial creativity (Fig 5, Fig 6). This means that, even though the seeds we provide are arbitrarily paired with the training datapoint, the models have learned to transform the arbitrary noise into meaningful randomness. Next, increasing the seed lengths consistently boosts algorithmic creativity for both next-token and multi-token approaches (Fig 523), presumably up to a limit. Thus, for Transformers, we propose viewing seed-conditioning as a knob for diversity distinct from temperature sampling.
Next, seed-conditioning results in algorithmic creativity comparable to temperature sampling. For the larger models (Fig 5), seed-conditioning generally outperforms the baseline “null” conditioning (which appears mode-collapsed, see Fig 181920). For the smaller model as well (Fig 6), we find gains in all tasks except for Sibling Discovery. Besides, for any fixed temperature, prefixing a seed only improves performance over a null prefix (Fig 618).
@@
4.2 An exploration of real-world summarization
For a more realistic examination, we finetune GPT models NTP and the multi-token teacherless objectives on summarization tasks (XSUM, CNN/DailyMail). We measure the diversity of a model for any given prompt by generating $5$ different completions and computing a Self-Bleu metric (Zhu et al., 2018). Admittedly though, a summarization task is not as open-ended as we would like: a higher quality model (i.e., higher Rouge; Lin, 2004) necessarily means lower diversity. To account for this, we plot how diversity evolves over time as a function of generation quality; we then find in Fig 7 that for a given model quality, the larger multi-token models achieve higher diversity (albeit only by a slight amount). This increase does not hold for smaller models and is not always noticeable for CNN/DailyMail (see §I). Interestingly, teacherless training consistently shows an increase in summarization quality, measured by Rouge.
@@ 4.2 An exploration of real-world summarization
Figure 7: Multi-token training improves diversity scores for XSUM summarization for large GPT-2 models: We observe gaps in diversity for a fixed quality over the course of finetuning.
@@ 5.1 Intuition for seed-conditioning
Seed-conditioning produces non-trivial algorithmic creativity even with greedy decoding. Perhaps, one could view the seed prefixes as a simpler alternative to varying the wordings of a prompt (Li et al., 2023; Lau et al., 2024; Naik et al., 2024; Li et al., 2022) or tuning a soft-prompt (Wang et al., 2024c), both of which are known to induce diversity. However, it is unexpected that the seeds provide any meaningful randomness at all, given they are arbitrarily picked with no semantic relationship to the training point. We do not understand this presently.
@@ 7 Related Work
Most related to seed-conditioning is DeSalvo et al. (2024) who induce diversity by varying a soft-prompt learned using a reconstruction loss. Our approach requires no modification to the architecture or the loss; however, we train the whole model, which is more expensive. A concurrent position paper (Jahrens & Martinetz, 2025) conceptually suggests injecting noise with the same motivation as us. Seed-conditioning may also be viewed as a controllable way to vary prompt wordings which is known to induce diverse outputs (Li et al., 2023; Lau et al., 2024; Naik et al., 2024; Li et al., 2022).
@@ 8 Conclusions
This work provides a minimal test-bed of tasks abstracting distinct modes of creativity. While these tasks are admittedly an extreme caricaturization of real-world tasks, they enable us to quantify otherwise elusive metrics like originality and diversity. They also enable us to control and investigate distinct parts of the current apparatus for language modeling (next-token learning and temperature sampling) and advocate for alternatives (multi-token learning and seed-conditioning). Next, one can investigate the effect of length generalization, scaling and the effect of in-context learning vs. finetuning. The surprising effectiveness of seed-conditioning raises various open questions (§B.1). Other profound questions arise, such as whether reasoning-enhancing methods like RL and CoT are optimal for enhancing open-ended diversity and originality (§5.2). We hope our work inspires discussion in the various directions of multi-token prediction, creativity and planning.
@@ Appendix A Transformer Training Objectives
[…]
@@ B.3 Examples of Triangle Discovery
    Antanaclasis: An antanaclasis involves using a word in two different senses in a sentence, while still ensuring that each sense has a coherent relationship with the rest of the sentence. Consider Benjamin Franklin’s quote, Your argument is sound, nothing but sound. Here, the two senses are $\texttt{sound}_{1}$ as in “logically correct”, $\texttt{sound}_{2}$ as in “noise”. This sentence encodes an pairwise relationship between three entities $\{\texttt{argument},\texttt{sound}_{1}$$\texttt{sound}_{2}\}$ individually. While the last two entities (the two senses) themselves must be related to each other (through the common word, sound), for a coherent sentence, both senses must also be appropriate descriptors for the first entity, argument. Thus, constructing this sentence requires searching through one’s vocabulary to discover three words that satisfy these three relationships simultaneously.
@@
      Devising “& Lit.” clues in cryptic crosswords are an altogether different, yet compelling example that require discovering a satisfying triangular relationship. Consider the clue “Some assassin in Japan” whose answer is Ninja. Here the phrase Some assassin in Japan participates in two senses. First, is the direct semantic sense as a definition of what a Ninja is. But there is a second, indirect sense: the word Some indicates that the solution lies as some substring of the phrase, namely ‘‘assassi(n in Ja)pan’’. Thus, constructing the clue requires identifying a triangular relationship between $\{\texttt{Ninja},\texttt{(Some assassin in Japan)}_{1},\texttt{(Some assassin in Japan)}_{2}\}$ just like in an antanaclasis. This is true generally of any & Lit. clues as these clues must perform “double duty” in pointing to the answer.
@@ C.1 Datasets inspired by combinational creativity
Next, we want a training dataset such that (a) the model can infer all the edges from the graph and yet (b) not all triangles appear in the dataset. This necessitates training on a dataset that consists not only of a subset of the triangles, but also of edges from the graph. Our training data consists of two parts: (1) $1/3$ are random triangles from the graph, (2) $2/3$ are random edges from the graph. In the training set, the triangle and edge samples are distinguished by a prefix “triangle:” or “edge:”. During test-time, we ensure that the model is prompted with “triangle:”. A triangle $(u,v,w)$ is tokenized as “tri: (u,v),(v,w),(w,u)(u,v),(v,w),(w,u)” and an edge $(u,v)$ as “edge: (u,v),(v,u)(u,v),(v,u)”. We provide both the directions of edge to potentially avoid any issues with the reversal curse (Berglund et al., 2024; Allen-Zhu & Li, 2023a).
For the default setting of the dataset, we set |V|=999|{V}|=999, $\texttt{deg}=3$, $\texttt{tri}=6$.
@@ Appendix D Further experimental details
Details for Gemma v1 (2B) model.  In Table 1, we provide the hyperparameter details for each of our datasets. We note some common details here. First, the batch size is $4$, but each sequence is packed with multiple examples; thus the model sequence length (divided by the input length) can be treated as a multiplicative factor that determines the effective batch size. The learning rates are chosen favorable to next-token prediction (not multi-token prediction). The training steps were chosen roughly based on a point after which the model had saturated in algorithmic creativity (and exhibited decreasing creativity). We use a learning rate with linear warm up for $100$ steps, followed by cosine annealing upto a factor $0.01\times$ of the maximum learning rate. To measure creativity, we sample a test dataset $T$ of $1024$ datapoints.
We represent the main tokens in our tasks with integers (ranging upwards of 0 to as many distinct integers as required). In the seed-conditioning setting, we use seeds of default length $10$, using randomly sampled uppercase characters from the English alphabet. In all datasets, we space-separate the vertices in a string, and comma-separate the edges.
@@
Hyperparameter
Sibling
Discovery
Triangle
Discovery
Circle
Construction
Line
Construction
@@
Weight given to
multi-token obj.
$0.5$ $0.5$ $0.75$ $0.75$
@@
Hyperparameter
Sibling
Discovery
Triangle
Discovery
Circle
Construction
Line
Construction
Max. Learning Rate $\times 10^{-4} 1×1$$\times 10^{-4}$ $\times 10^{-4} 1×1$$\times 10^{-4}$
@@
Weight given to
multi-token obj.
$0.5$ $0.33$ $0.75$ $0.75$
@@ E.1 Varying sampling methods
Fig 18, Fig 19, and Fig 20 report creativity, memorization and coherence (i.e., fraction of generated strings that are coherent) for various sampling methods (greedy decoding and nucleus sampling) with various prefix conditionings. Here we not only look at the baseline (denoted by null) and seed prefix (denoted by seed), we also look at using a prefix of pause tokens (Goyal et al., 2024) (denoted by pause). This can be viewed as way to provide extra computation to the model before it emits its outputs. If this had the same effect of seed-conditioning, then we would conclude that the gains of seed-conditioning come purely from prepending extra computation to the model — not from the act of injecting noise at the input. However, in our experiments, we do not find noticeable gains from pause-conditioning, establishing that noise-injection is in itself beneficial.